Determine the direction of maximum increase of the scalar field f(x,y,z)=xe^y+z^2 at the point O( 1,ln2,3).


1
Expert's answer
2021-01-20T10:28:59-0500

Answer

Gradient of g at P0

"\nabla g|_{P_0}=\begin{pmatrix}
 e^y\\
 xe^y\\
2z
\end{pmatrix}_{(1, ln2, 3) }"

"=\begin{pmatrix}
 2\\
 2\\
2/3
\end{pmatrix}"

The direction in which g increase most rapidly

"u=\frac{3}{\sqrt{19}}\begin{pmatrix}
 1 \\
1\\
 1/3
\end{pmatrix}"